# Assisting Mathematical Formalization with A Learning-based Premise Retriever

### v2

@@ 4.4.1 Effect of Data Perturbation
Considering the generally high quality of proofs in Mathlib, models trained on this dataset may be sensitive to the common issues of redundant variables, missing conditions, and disordered sequences in the context of proof states, resulting in poor generalization ability. We evaluate our retrieval model’s robustness to low-quality inputs by perturbing the query states in the test set. Specifically, we applied two perturbation strategies to a subset of the test data: shuffling the context or randomly removing 20% of the context from states with a context length of 15 or more. As shown in Figure 5, all the metrics exhibit a moderate decline as the perturbation ratio increases, with a maximum drop of approximately 6% when k=5,10 compared to the unperturbed case. Especially, the model’s performance even improves when k=1. These results demonstrate that our retrieval model is robust to variations in the order of query states and the potential absence of certain local hypotheses.
4.4.2 Impact of Reduced Training Data
@@
The sparsity of data is an inherent challenge in formalization tasks. We test the sensitivity of our retrieval model to the amount of training data. The tests are conducted on the Random split dataset. The results in Figures 5 show that the model trained on the full dataset outperforms the model trained on only 25% of the data by approximately 25% in terms of k=5,10. The consistent performance improvement suggests that there is still potential for further enhancement by increasing the amount of training data. When k=1, the performance seems to reach a bottleneck as the data size increases. Therefore, it may be necessary to incorporate re-ranking or explore alternative methods to further improve performance.
@@
4.5 Theorem Proving Experiments (RQ4)
As mentioned in Section 3.3.3, we fine-tune ByT5 [40] and obtain a tactic generator. Facilitating the generator, we conduct theorem-proving experiments with retrieval. The results in Figure 6 show that assisted by our model, the generator averagely performs better than ReProver [41]. Additionally, we evaluate our method on the Minif2f benchmark [45] using the model trained on the random split dataset. Our approach achieves a pass@1 rate of 30.74%, while ReProver reached 28.28%, indicating that our method is more effective. This means that retrieval-augmented language models are effective: For tasks such as formal theorem proving, which relies on external theorem libraries, an effective retrieval model is expected to enhance generation performance.
On test sets with high premise frequency but relatively short proof, our model performs better than ReProver, demonstrating the advantages of a more powerful retriever. In contrast, on test sets with long proof but low premise frequency, our model performs almost on par with ReProver. However, on the RD test sets, our model performs slightly worse than ReProver, despite the improvement in retrieval accuracy. We attribute this outcome to several factors. First, incorporating retrieved premises into the prompt increases the length of the context that the model must process compared to providing only the proof state. This imposes certain requirements on the model’s parameter size and architecture, as smaller models may struggle to accurately identify useful information within longer contexts. We will explore this issue in our future work. Additionally, since Lean users are continuously developing new automated tactics, solving simpler problems often depends more on selecting the appropriate tactic than on searching for relevant theorems, making tactic selection a more efficient approach in such cases.
@@ 5 Conclusion
In this paper, we propose an innovative method that leverages data extracted from Mathlib4 to train a retrieval model specifically designed for Lean premise retrieval. Experiments show that our method outperforms previous models and demonstrates the potential of domain-specific smaller models in data-sparse tasks, such as formal premise retrieval. As shown in Figure 7, we have developed a search engine for premise retrieval based on our model. We hope that our work can help accelerate mathematical formalization and contributes to the researchers in the community.
Figure 7: The interface of our premise retrieval engine.
Although our method has yielded promising results, there is still room for optimization in the retriever’s backbone model. For the Lean premise retrieval task, the data corpus contains valuable semantic information, such as the hierarchy of Lean’s type system, that has yet to be fully explored. In the future, we will focus on leveraging this untapped semantic information more effectively and developing better formal language models.
Acknowledgments
Yicheng Tao and Shanwen Wang are supported by National Key Technologies R & D Program of China, Grant №: 2024YFA1014001. We thank Beijing International Center for Mathematical Research(BICMR) at Peking University and Institute for Mathematical Sciences(IMS) at National University of Singapore for organizing Lean workshops, which provides great opportunities for learning and communication.

### v3

@@ 4.4 Robustness Experiments
Effect of Data Perturbation Considering the generally high quality of proofs in Mathlib, models trained on this dataset may be sensitive to the common issues of redundant variables, missing conditions, and disordered sequences in the context of proof states, resulting in poor generalization ability. We evaluate our retrieval model’s robustness to low-quality inputs by perturbing the query states in the test set. Specifically, we applied two perturbation strategies to a subset of the test data: shuffling the context or randomly removing 20% of the context from states with a context length of 15 or more. As shown in Figure 6, all the metrics exhibit a moderate decline as the perturbation ratio increases, with a maximum drop of approximately 6% when $k=5,10$ compared to the unperturbed case. Especially, the model’s performance even improves when $k=1$. These results demonstrate that our retrieval model is robust to variations in the order of query states and the potential absence of certain local hypotheses.
@@
Impact of Reduced Training Data The sparsity of data is an inherent challenge in formalization tasks. We test the sensitivity of our retrieval model to the amount of training data. The tests are conducted on the Random split dataset. The results in Figures 6 show that the model trained on the full dataset outperforms the model trained on only 25% of the data by approximately 25% in terms of $k=5,10$. The consistent performance improvement suggests that there is still potential for further enhancement by increasing the amount of training data. When $k=1$, the performance seems to reach a bottleneck as the data size increases. Therefore, it may be necessary to incorporate re-ranking or explore alternative methods to further improve performance.
Figure 6: The variation of the normalized values of the model’s metrics under different data perturbations and different training data proportion. In data perturbation experiments, the metric with no perturbation $v_{0}$ is adopted as baseline. The normalized value $v^{*}=v/v_{0}$. In data proportion experiments, the metric with 25% training data is adopted as baseline.
4.5 Theorem Proving Experiments
As mentioned in Section 3.3, we fine-tune ByT5 (Xue et al., 2022) and obtain a tactic generator. Facilitating the generator, we conduct theorem-proving experiments with retrieval. The results in Figure 7 show that assisted by our model, the generator averagely performs better than ReProver (Yang et al., 2024). Additionally, we evaluate our method on the Minif2f benchmark (Zheng et al., 2022) using the model trained on the random split dataset. Our approach achieves a pass@1 rate of 30.74%, while ReProver reached 28.28%, indicating that our method is more effective. This means that retrieval-augmented language models are effective: For tasks such as formal theorem proving, which relies on external theorem libraries, an effective retrieval model is expected to enhance generation performance.
On test sets with high premise frequency but relatively short proof, our model performs better than ReProver, demonstrating the advantages of a more powerful retriever. In contrast, on test sets with long proof but low premise frequency, our model performs almost on par with ReProver. However, on the RD test sets, our model performs slightly worse than ReProver, despite the improvement in retrieval accuracy. We attribute this outcome to several factors. First, incorporating retrieved premises into the prompt increases the length of the context that the model must process compared to providing only the proof state. This imposes certain requirements on the model’s parameter size and architecture, as smaller models may struggle to accurately identify useful information within longer contexts. We will explore this issue in our future work.
@@ 5 Conclusion
In this paper, we propose an innovative method that leverages data extracted from Mathlib4 to train a retrieval model specifically designed for Lean premise retrieval. Experiments show that our method outperforms previous models and demonstrates the potential of domain-specific smaller models in data-sparse tasks, such as formal premise retrieval. We have developed a search engine for premise retrieval based on our model. We hope that our work can help accelerate mathematical formalization and contributes to the researchers in the community.
Although our method has yielded promising results, there is still room for optimization in the retriever’s backbone model. For the Lean premise retrieval task, the data corpus contains valuable semantic information, such as the hierarchy of Lean’s type system, that has yet to be fully explored. Moreover, the results of theorem proving experiments need further clarification with stronger provers, such as DeepSeek-Prover-V2 (Ren et al., 2025), to demonstrate the effects of premise selection for automated theorem proving. In the future, we will focus on leveraging this untapped semantic information more effectively and developing better models.
Acknowledgements
Yicheng Tao and Shanwen Wang are supported by National Key Technologies R&D Program of China, Grant No. 2024YFA1014001. Haotian Liu and Hongteng Xu are supported by Beijing Natural Science Foundation (Grant No. L233008), the fund for building world-class universities of Renmin University of China, and the funds from Beijing Key Laboratory of Research on Large Models and Intelligent Governance, and Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education. We thank Beijing International Center for Mathematical Research(BICMR) at Peking University and Institute for Mathematical Sciences(IMS) at National University of Singapore for organizing Lean workshops, which provides great opportunities for learning and communication.
Impact Statement
The formalization community has been exploring the potential of AI tools as assistants. In this work we develop an intelligent search engine for formal theorems based on proof states, aiming to accelerating the process of mathematical formalization.
Appendix A Data Extraction
Several tools have been developed to extract data from a Lean project. To meet our requirements, we utilize the script from LeanDojo and incorporate additional features. We applied this pipeline to Mathlib [^4] We used tag v4.10.0 of mathlib4. https://github.com/leanprover-community/mathlib4/tree/v4.10.0. The extracted information covers all the premises and proofs in this project, along with metadata such as file dependencies. For each premise, we extract its argument list and output. The proofs are lists of tactics, with the state before/after the tactics and premises used in each step. Figure 8 illustrates the statistics related to context length and the number of premises in the dataset. The context length represents the length of $\Gamma$, while the premise number indicates how many premises are utilized in a specific tactic. It is important to note that tactics that do not have any premises are excluded from these statistics.